### v2

@@ 3.2 Large Kernel Convolution Attention (LKCA)
Large Kernel Convolutional Attention (LKCA) can be expressed equivalently as the attention operation with shared parameters described above in terms of code and mathematical implementation. Listing 1.2 showcases pseudocode for implementing LKCA from the perspective of convolution. The convolutional kernel is a learnable weight matrix of size $N^{\frac{1}{2}}\times 2-1$, and the input feature map is a matrix $V$ of size $N^{\frac{1}{2}}$ with a channel count equal to the hidden dimension. After zero-padding of size $N^{\frac{1}{2}}-1$, the convolution operation is applied to the two matrices.
@@
class Attention_CNN(nn.Module):
    def __init__(num_patches):
        super().__init__()
        n = num_patches
        H, W = n**0.5
        weight = nn.Parameter(torch.zeros(H*2-1, W*2-1))
    def forward(X):
        b, n, d = X.shape
        V= nn.Linear(X)
        V = rearrange(V, ’b (H W) d -> (b d) H W’)
        Out = F.conv2d(V, weight, zero_padding=(H-1, W-1))
        Out = rearrange(Out, ’(b d) H W -> b (H W) d’)
        return Out
Listing 2: Pseudocode from convolution perspective
3.2.3 Overall Architecture.
@@
$\mathbf{z}_{0}$ $=\left[\mathbf{x}_{p}^{1}\mathbf{E};\mathbf{x}_{p}^{2}\mathbf{E};\cdots;\mathbf{x}_{p}^{N}\mathbf{E}\right]+\mathbf{E}_{pos},$ $\mathbf{E}\in\mathbb{R}^{\left(P^{2}\cdot C\right)\times D},\mathbf{E}_{pos}\in\mathbb{R}^{N\times D}$ (1)
$\mathbf{z}_{\ell}^{\prime}$ $=\operatorname{LKCA}\left(\operatorname{LN}\left(\mathbf{z}_{\ell-1}\right)\right)+\mathbf{z}_{\ell-1},$ $\ell=1\ldots L$ (2)
$\mathbf{z}_{\ell}$ $=\operatorname{MLP}\left(\operatorname{LN}\left(\mathbf{z}_{\ell}^{\prime}\right)\right)+\mathbf{z}_{\ell}^{\prime},$ $\ell=1\ldots L$ (3)
@@
4 Experiments
4.1 Image Classification
@@
Experimental Setting. In our experimental setup, we conducted extensive image classification experiments on four prominent datasets: CIFAR-10, CIFAR-100, SVHN, and Tiny-ImageNet. All four datasets are characterized by low resolution, with image sizes of 32 for CIFAR-10, CIFAR-100, and SVHN, and 64 for Tiny-ImageNet. The training regime was consistently maintained across all experiments, conducted on a single RTX 2080ti GPU. The training spanned 100 epochs, employing 4 data loading workers, a batch size of 128. For optimizer, we choose AdamW [62]. The default learning rate was set to 0.001, with a weight decay of 0.05. We utilized cosine annealing with a warm-up period of 10 epochs. During the training phase, we employed Stochastic Depth [63], Label Smoothing [45], and Random Erasing [64] with a probability of 0.25, a maximum erasing area of 0.4, and an aspect of erasing area of 0.3 as regularization strategies to enhance the robustness and generalization capabilities of our model. Regarding data augmentation, we employed various techniques. Specifically, we applied RandomHorizontalFlip, RandomCrop [65], Mixup [66], and Cutmix [67]. Additionally, we incorporated Autoaugmentation [68], utilizing the CIFAR10 Policy for CIFAR-10 and CIFAR-100, the SVHN Policy for SVHN, and the ImageNet Policy for Tiny-ImageNet.
@@
Comparison of ViT and LKCA. Initially, we conducted extensive experiments on CIFAR-10, CIFAR-100, and Tiny-ImageNet to compare the performance of the original ViT with an LKCA-modified version of ViT having a similar order of magnitude in terms of parameters. The purpose of these experiments was to validate the effectiveness of LKCA across various parameter magnitudes and datasets. The parameter sizes of both the ViT and the LKCA-modified ViT varied across a range of magnitudes, spanning from small-scale models with 0.5 million parameters to large-scale models with 10 million parameters. Specifically, the parameter magnitudes were explored at levels of 0.5M, 1M, 2M, 4M, and 8M. It is important to note that the benchmark representing the model size across all variants is fixed at 2.69 million parameters, and it is independent of any subjective categorization as either large or small. This standardized model size serves as the reference point for assessing the performance of the different variants, ensuring a consistent comparison across the experiment.
@@
Table 1: Performance Comparison of Vision Transformer and LKCA Variants on Image Classification Tasks on CIFAR-10 (Top-1 accuracy)
Model Acc # Params Flops
ViT 93.36 2.69M 174.25M
ViT-LKCA 94.11 (+0.75) 2.69M 170.81M
VIT/0.5M 87.83 0.48M 30.58M
ViT/0.5M-LKCA 90.94 (+3.11) 0.50M 31.34M
VIT/1M 91.16 1.06M 68.36M
ViT/1M-LKCA 92.49 (+1.33) 1.03M 64.82M
VIT/2M 92.32 1.96M 126.09M
ViT/2M-LKCA 93.55 (+1.23) 2.03M 128.66M
VIT/4M 93.64 3.95M 255.47M
ViT/4M-LKCA 94.71 (+1.07) 4.05M 257.43M
VIT/8M 94.46 7.85M 508.72M
ViT/8M-LKCA 94.88 (+0.42) 7.98M 506.00M
Table 1 presents a performance comparison of ViT and its LKCA variants on the CIFAR-10 image classification task, along with their performance across different orders of magnitude in terms of parameters. The table includes the Top-1 accuracy (Acc), the number of parameters (# Params), and Floating Point Operations (Flops) for each model. ViT and ViT-LKCA: The ViT model achieves a Top-1 accuracy of 93.36%, with 2.69M parameters and 174.25M Flops. The ViT-LKCA outperforms with a Top-1 accuracy of 94.11% (an improvement of 0.75 percentage points) and has 2.69M parameters and 170.81M Flops. Small-scale model (0.5M parameter magnitude): The ViT/0.5M model has a Top-1 accuracy of 87.83%, 0.48M parameters, and 30.58M Flops. The ViT/0.5M-LKCA, with LKCA improvement, significantly improves accuracy to 90.94% (an increase of 3.11 percentage points) with 0.50M parameters and 31.34M Flops. Medium-scale models (1M, 2M, 4M parameter magnitudes): For ViT and ViT-LKCA models with 1M, 2M, and 4M parameters, LKCA consistently yields higher Top-1 accuracy. For instance, ViT/1M-LKCA achieves a 1.33 percentage point improvement, reaching an accuracy of 92.49%. Large-scale model (8M parameter magnitude): For ViT and ViT-LKCA models with 8M parameters, the performance gap is narrower. ViT/8M-LKCA slightly improves Top-1 accuracy by 0.42 percentage points, reaching 94.88%.
@@
Table 2: Performance Comparison of Vision Transformer and LKCA Variants on Image Classification Tasks on CIFAR-100 (Top-1 accuracy)
Model Acc # Params Flops
ViT 72.65 2.69M 174.25M
ViT-LKCA 76.50 (+3.85) 2.69M 170.81M
VIT/0.5M 64.68 0.48M 30.58M
ViT/0.5M-LKCA 69.35 (+4.67) 0.50M 31.34M
VIT/1M 70.86 1.06M 68.36M
ViT/1M-LKCA 73.40 (+2.54) 1.03M 64.82M
VIT/2M 71.17 1.96M 126.09M
ViT/2M-LKCA 75.94 (+4.77) 2.03M 128.66M
VIT/4M 72.82 3.95M 255.47M
ViT/4M-LKCA 76.95 (+4.13) 4.05M 257.43M
VIT/8M 73.39 7.85M 508.72M
ViT/8M-LKCA 78.57 (+5.18) 7.98M 506.00M
Table 2 illustrates Acomparison between ViT and LKCA in the context of the CIFAR-100. ViT and ViT-LKCA: The base ViT model achieves a Top-1 accuracy of 72.65% with 2.69M parameters and 174.25M Flops. In contrast, ViT-LKCA exhibits superior performance, attaining a Top-1 accuracy of 76.50% (an increase of 3.85 percentage points) with the same 2.69M parameters and 170.81M Flops. Small-scale model (0.5M parameter magnitude): The ViT/0.5M model achieves a Top-1 accuracy of 64.68% with 0.48M parameters and 30.58M Flops. The ViT/0.5M-LKCA, featuring LKCA enhancements, significantly improves accuracy to 69.35% (a notable increase of 4.67 percentage points) with 0.50M parameters and 31.34M Flops. Medium-scale models (1M, 2M, 4M parameter magnitudes): ViT and ViT-LKCA models with 1M, 2M, and 4M parameters consistently benefit from LKCA improvements. For instance, ViT/2M-LKCA achieves a noteworthy 4.77 percentage point improvement and reaches an accuracy of 75.94%. Large-scale model (8M parameter magnitude): For ViT and ViT-LKCA models with 8M parameters, the performance difference is more subtle. ViT/8M-LKCA exhibits a huge improvement of 5.18 percentage points, reaching a Top-1 accuracy of 78.57%.
Table 3: Performance Comparison of Vision Transformer and LKCA Variants on Image Classification Tasks on Tiny-ImageNet (Top-1 accuracy)
Model Acc # Params Flops
ViT 55.74 2.69M 174.25M
ViT-LKCA 60.95 (+5.21) 2.69M 170.81M
VIT/0.5M 46.53 0.48M 30.58M
ViT/0.5M-LKCA 50.59 (+4.06) 0.50M 31.34M
VIT/1M 53.46 1.06M 68.36M
ViT/1M-LKCA 57.29 (+3.83) 1.03M 64.82M
VIT/2M 54.58 1.96M 126.09M
ViT/2M-LKCA 60.70 (+6.12) 2.03M 128.66M
VIT/4M 56.16 3.95M 255.47M
ViT/4M-LKCA 62.30 (+6.14) 4.05M 257.43M
VIT/8M 57.39 7.85M 508.72M
ViT/8M-LKCA 63.38 (+5.99) 7.98M 506.00M
@@
Table 3 provides the comparison on Tiny-ImageNet. ViT and ViT-LKCA: The baseline ViT model achieves a Top-1 accuracy of 55.74% with 2.69M parameters and 174.25M Flops. In contrast, the ViT-LKCA model outperforms significantly with a Top-1 accuracy of 60.95% (an increase of 5.21 percentage points) and has the same 2.69M parameters and 170.81M Flops. Small-scale model (0.5M parameter magnitude): The ViT/0.5M model achieves a Top-1 accuracy of 46.53% with 0.48M parameters and 30.58M Flops. The ViT/0.5M-LKCA, incorporating LKCA enhancements, significantly improves accuracy to 50.59% (an increase of 4.06 percentage points) with 0.50M parameters and 31.34M Flops. Medium-scale models (1M, 2M, 4M parameter magnitudes): With ViT/4M-LKCA achieving a remarkable 6.14 percentage point improvement and reaching an accuracy of 62.30%. Large-scale model (8M parameter magnitude): For ViT and ViT-LKCA models with 8M parameters, ViT/8M-LKCA exhibits a significant improvement of 5.99 percentage points, achieving a Top-1 accuracy of 63.38%.
@@
Comparison of ViT Variants and LKCA. Next, we compared LKCA with other mainstream variants of Vision Transformer.
@@
Table 4: Comparison of LKCA with Several Mainstream ViT Variants on Different Datasets (Top-1 accuracy)
Model CIFAR10 CIFAR100 SVHN # Params Flops
ViT-Lite [3] 91.16 70.86 97.53 1.06M 68.36M
Swin-T [53] 91.21 70.37 96.40 1.01M 38.26M
CaiT-T [69] 91.99 70.65 97.53 1.00M 61.16M
MobileViTv2-T [70] 91.46 67.88 96.60 1.00M 7.43M
LKCA-T 92.49 (Ours) 73.40 (Ours) 97.15 (Ours) 1.03M 64.82M
ViT-Small [71] 93.36 72.65 97.90 2.69M 174.25M
Swin-S [53] 93.99 75.34 97.59 2.87M 94.72M
CaiT-S [69] 94.02 74.70 97.79 2.74M 163.55M
MobileViTv2-S [70] 93.48 71.70 96.98 3.92M 24.55M
LKCA-B 94.11 (Ours) 76.50 (Ours) 97.68 (Ours) 2.69M 170.81M
From the information in the table 4. In the context of small-scale models, Swin-T and CaiT-T perform similarly in terms of Top-1 accuracy, while MobileViTv2 lags slightly behind but possesses a significant advantage in computational efficiency. Compared to other small-scale models, LKCA-T achieves a higher Top-1 accuracy. Moving to large-scale models, it is a common trend that they generally outperform their small-scale counterparts. LKCA-B stands out by significantly surpassing ViT and MobileViTv2 in Top-1 accuracy, and it slightly outperforms CaiT-S and Swin-S. Overall, in CIFAR10 and CIFAR100, LKCA demonstrates higher Top-1 accuracy compared to other models. However, there is no clear advantage observed in the SVHN dataset.
Finally, we have validated the performance of various ViT variants and the LKCA model on the Tiny-ImageNet dataset, while also providing details on the models’ parameter counts and computational complexities in Table 5.
@@
Table 5: Comparison of Various ViT Variants with LKCA on Tiny-ImageNet Dataset (Top-1 accuracy)
Model Tiny-ImageNet # Params Flops
T2T-T [72] 53.92 1.07M 78.23M
RvT-T [73] 50.65 1.10M 57.61M
Swin-T [53] 54.93 1.06M 38.90M
CaiT-T [69] 54.76 1.03M 61.85M
XCiT-T [74] 56.78 0.96M 51.44M
ViT-Lite [3] 53.46 1.11M 69.64M
DeepViT-T [60] 34.64 0.99M 62.96M
RegionViT-T [59] 54.32 0.97M 29.38M
CrossViT-T [75] 47.03 1.04M 57.59M
LKCA-T 57.29 (Ours) 1.07M 66.10M
T2T-S [72] 41.25 2.56M 52.96M
RvT-S [73] 55.51 2.72M 145.09M
Swin-S [53] 58.61 2.93M 95.55M
CaiT-S [69] 59.21 2.77M 164.46M
XCiT-S [74] 60.09 2.81M 157.54M
ViT-Small [3] 55.74 2.76M 176.06M
DeepViT-S [60] 44.45 2.54M 162.85M
Twins_SVT-S [55] 37.13 2.76M 197.00M
RegionViT-S [59] 53.96 2.86M 53.82M
CrossViT-S [75] 52.70 2.40M 126.11M
LKCA-B 60.95 (Ours) 2.76M 172.78M
T2T-B [72] 58.46 13.45M 853.02M
CvT-B [71] 55.88 6.52M 102.56M
MobileViTv2 [70] 58.28 8.17M 189.77M
Twins_SVT-B [55] 49.24 9.04M 308.74M
RegionViT-B [59] 57.83 12.39M 195.02M
LKCA-L 63.43 (Ours) 12.65M 802.99M
4.2 Semantic Segmentation

### v3

@@ III-B2 From the Perspective of Convolution: Implementing LKCA
Large Kernel Convolutional Attention (LKCA) can be expressed equivalently as the attention operation with shared parameters described above in terms of code and mathematical implementation. Algorithm 2 showcases pseudocode for implementing LKCA from the perspective of convolution. The convolutional kernel is a learnable weight matrix of size $N^{\frac{1}{2}}\times 2-1$, and the input feature map is a matrix $V$ of size $N^{\frac{1}{2}}$ with a channel count equal to the hidden dimension. After zero-padding of size $N^{\frac{1}{2}}-1$, the convolution operation is applied to the two matrices.
@@
Algorithm 2 Forward propagation process of Attention-CNN
1:function Attention_CNN_Forward($X$)
2:     input: $X$ – input tokens, where $X\in\mathbb{R}^{b\times n\times d}$
3:     initialize: $H,W\leftarrow\sqrt{n}$, $weight\leftarrow$ zero tensor of shape $(H*2-1,W*2-1)$
4:     $V\leftarrow\text{Linear}(X)$
5:     $V\leftarrow\text{rearrange}(V,^{\prime}b(HW)d\rightarrow(bd)HW^{\prime})$
6:     $Out\leftarrow\text{conv2d}(V,weight,\text{zero\_padding}=(H-1,W-1))$
7:     $Out\leftarrow\text{rearrange}(Out,^{\prime}(bd)HW\rightarrow b(HW)d^{\prime})$
8:     return $Out$
9:end function
III-B3 Overall Architecture
@@
$\mathbf{z}_{0}$ $=\left[\mathbf{x}_{p}^{1}\mathbf{E};\mathbf{x}_{p}^{2}\mathbf{E};\cdots;\mathbf{x}_{p}^{N}\mathbf{E}\right]+\mathbf{E}_{pos},$ (1)
$\quad\mathbf{E}\in\mathbb{R}^{\left(P^{2}\cdot C\right)\times D},\quad\mathbf{E}_{pos}\in\mathbb{R}^{N\times D}$ (2)
$\mathbf{z}_{\ell}^{\prime}$ $=\operatorname{LKCA}\left(\operatorname{LN}\left(\mathbf{z}_{\ell-1}\right)\right)+\mathbf{z}_{\ell-1},$ $\quad\ell=1\ldots L$ (3)
$\mathbf{z}_{\ell}$ $=\operatorname{MLP}\left(\operatorname{LN}\left(\mathbf{z}_{\ell}^{\prime}\right)\right)+\mathbf{z}_{\ell}^{\prime},$ $\quad\ell=1\ldots L$ (4)
@@
TABLE I: Comparison of LKCA with Several Mainstream ViT Variants on Different Datasets (Top-1 accuracy)
Model CIFAR10 CIFAR100 SVHN # Params
ViT-Lite [1] 91.16 70.86 97.53 1.06M
Swin-T [58] 91.21 70.37 96.40 1.01M
CaiT-T [67] 91.99 70.65 97.53 1.00M
MobileViTv2-T [68] 91.46 67.88 96.60 1.00M
LKCA-T 92.49 (Ours) 73.40 (Ours) 97.15 (Ours) 1.03M
ViT-Small [69] 93.36 72.65 97.90 2.69M
Swin-S [58] 93.99 75.34 97.59 2.87M
CaiT-S [67] 94.02 74.70 97.79 2.74M
MobileViTv2-S [68] 93.48 71.70 96.98 3.92M
LKCA-B 94.11 (Ours) 76.50 (Ours) 97.68 (Ours) 2.69M
IV Experiments
IV-A Experimental Setup
@@
In our experimental setup, we conduct extensive image classification experiments on four prominent datasets: CIFAR-10, CIFAR-100, SVHN, and Tiny-ImageNet. All four datasets are characterized by low resolution, with image sizes of 32 for CIFAR-10, CIFAR-100, and SVHN, and 64 for Tiny-ImageNet. The training regime is consistently maintained across all experiments, conducted on a single RTX 2080ti GPU. The training spans 100 epochs, employing 4 data loading workers, and a batch size of 128. For optimizer, we choose AdamW [70]. The default learning rate is set to 0.001, with a weight decay of 0.05. We utilize cosine annealing with a warm-up period of 10 epochs. During the training phase, we employ Stochastic Depth [71], Label Smoothing [50], and Random Erasing [72] with a probability of 0.25, a maximum erasing area of 0.4, and an aspect of erasing area of 0.3 as regularization strategies to enhance the robustness and generalization capabilities of our model. Regarding data augmentation, we employ various techniques. Specifically, we apply RandomHorizontalFlip, RandomCrop [73], Mixup [74], and Cutmix [75]. Additionally, we incorporate Autoaugmentation [76], utilizing the CIFAR10 Policy for CIFAR-10 and CIFAR-100, the SVHN Policy for SVHN, and the ImageNet Policy for Tiny-ImageNet.
IV-B Comparison of ViT and LKCA.
@@
Initially, we conduct extensive experiments on CIFAR-10, CIFAR-100, and Tiny-ImageNet to compare the performance of the original ViT with an LKCA-modified version of ViT having a similar order of magnitude in terms of parameters. The purpose of these experiments is to validate the effectiveness of LKCA across various parameter magnitudes and datasets. The parameter sizes of both the ViT and the LKCA-modified ViT vary across a range of magnitudes, spanning from small-scale models with 0.5 million parameters to large-scale models with 10 million parameters. Specifically, the parameter magnitudes are explored at levels of 0.5M, 1M, 2M, 4M, and 8M. It is important to note that the benchmark representing the model size across all variants is fixed at 2.69 million parameters, and it is independent of any subjective categorization as either large or small. This standardized model size serves as the reference point for assessing the performance of the different variants, ensuring a consistent comparison across the experiment.
@@
TABLE II: Performance Comparison of Vision Transformer and LKCA Variants on Image Classification Tasks on CIFAR-10 (Top-1 accuracy)
Model Acc # Params
ViT 93.36 2.69M
ViT-LKCA 94.11 (+0.75) 2.69M
VIT/0.5M 87.83 0.48M
ViT/0.5M-LKCA 90.94 (+3.11) 0.50M
VIT/1M 91.16 1.06M
ViT/1M-LKCA 92.49 (+1.33) 1.03M
VIT/2M 92.32 1.96M
ViT/2M-LKCA 93.55 (+1.23) 2.03M
VIT/4M 93.64 3.95M
ViT/4M-LKCA 94.71 (+1.07) 4.05M
VIT/8M 94.46 7.85M
ViT/8M-LKCA 94.88 (+0.42) 7.98M
@@
TABLE III: Performance Comparison of Vision Transformer and LKCA Variants on Image Classification Tasks on CIFAR-100 (Top-1 accuracy)
Model Acc # Params
ViT 72.65 2.69M
ViT-LKCA 76.50 (+3.85) 2.69M
VIT/0.5M 64.68 0.48M
ViT/0.5M-LKCA 69.35 (+4.67) 0.50M
VIT/1M 70.86 1.06M
ViT/1M-LKCA 73.40 (+2.54) 1.03M
VIT/2M 71.17 1.96M
ViT/2M-LKCA 75.94 (+4.77) 2.03M
VIT/4M 72.82 3.95M
ViT/4M-LKCA 76.95 (+4.13) 4.05M
VIT/8M 73.39 7.85M
ViT/8M-LKCA 78.57 (+5.18) 7.98M
@@
Table II presents a performance comparison of ViT and its LKCA variants on the CIFAR-10 image classification task, along with their performance across different orders of magnitude in terms of parameters. The table includes the Top-1 accuracy (Acc), the number of parameters (# Params), and Floating Point Operations (Flops) for each model. ViT and ViT-LKCA: The ViT model achieves a Top-1 accuracy of 93.36%, with 2.69M parameters and 174.25M Flops. The ViT-LKCA outperforms with a Top-1 accuracy of 94.11% (an improvement of 0.75 percentage points) and has 2.69M parameters and 170.81M Flops. Small-scale model (0.5M parameter magnitude): The ViT/0.5M model has a Top-1 accuracy of 87.83%, 0.48M parameters, and 30.58M Flops. The ViT/0.5M-LKCA, with LKCA improvement, significantly improves accuracy to 90.94% (an increase of 3.11 percentage points) with 0.50M parameters and 31.34M Flops. Medium-scale models (1M, 2M, 4M parameter magnitudes): For ViT and ViT-LKCA models with 1M, 2M, and 4M parameters, LKCA consistently yields higher Top-1 accuracy. For instance, ViT/1M-LKCA achieves a 1.33 percentage point improvement, reaching an accuracy of 92.49%. Large-scale model (8M parameter magnitude): For ViT and ViT-LKCA models with 8M parameters, the performance gap is narrower. ViT/8M-LKCA slightly improves Top-1 accuracy by 0.42 percentage points, reaching 94.88%.
Table III illustrates a comparison between ViT and LKCA in the context of the CIFAR-100. ViT and ViT-LKCA: The base ViT model achieves a Top-1 accuracy of 72.65% with 2.69M parameters and 174.25M Flops. In contrast, ViT-LKCA exhibits superior performance, attaining a Top-1 accuracy of 76.50% (an increase of 3.85 percentage points) with the same 2.69M parameters and 170.81M Flops. Small-scale model (0.5M parameter magnitude): The ViT/0.5M model achieves a Top-1 accuracy of 64.68% with 0.48M parameters and 30.58M Flops. The ViT/0.5M-LKCA, featuring LKCA enhancements, significantly improves accuracy to 69.35% (a notable increase of 4.67 percentage points) with 0.50M parameters and 31.34M Flops. Medium-scale models (1M, 2M, 4M parameter magnitudes): ViT and ViT-LKCA models with 1M, 2M, and 4M parameters consistently benefit from LKCA improvements. For instance, ViT/2M-LKCA achieves a noteworthy 4.77 percentage point improvement and reaches an accuracy of 75.94%. Large-scale model (8M parameter magnitude): For ViT and ViT-LKCA models with 8M parameters, the performance difference is more subtle. ViT/8M-LKCA exhibits a huge improvement of 5.18 percentage points, reaching a Top-1 accuracy of 78.57%.
TABLE IV: Performance Comparison of Vision Transformer and LKCA Variants on Image Classification Tasks on Tiny-ImageNet (Top-1 accuracy)
Model Acc # Params
ViT 55.74 2.69M
ViT-LKCA 60.95 (+5.21) 2.69M
VIT/0.5M 46.53 0.48M
ViT/0.5M-LKCA 50.59 (+4.06) 0.50M
VIT/1M 53.46 1.06M
ViT/1M-LKCA 57.29 (+3.83) 1.03M
VIT/2M 54.58 1.96M
ViT/2M-LKCA 60.70 (+6.12) 2.03M
VIT/4M 56.16 3.95M
ViT/4M-LKCA 62.30 (+6.14) 4.05M
VIT/8M 57.39 7.85M
ViT/8M-LKCA 63.38 (+5.99) 7.98M
TABLE V: Comparison of Various ViT Variants with LKCA on Tiny-ImageNet Dataset (Top-1 accuracy)
Model Tiny-ImageNet # Params
T2T-T [77] 53.92 1.07M
RvT-T [78] 50.65 1.10M
Swin-T [58] 54.93 1.06M
CaiT-T [67] 54.76 1.03M
XCiT-T [79] 56.78 0.96M
ViT-Lite [1] 53.46 1.11M
DeepViT-T [65] 34.64 0.99M
RegionViT-T [64] 54.32 0.97M
CrossViT-T [80] 47.03 1.04M
LKCA-T 57.29 (Ours) 1.07M
T2T-S [77] 41.25 2.56M
RvT-S [78] 55.51 2.72M
Swin-S [58] 58.61 2.93M
CaiT-S [67] 59.21 2.77M
XCiT-S [79] 60.09 2.81M
ViT-Small [1] 55.74 2.76M
DeepViT-S [65] 44.45 2.54M
Twins_SVT-S [60] 37.13 2.76M
RegionViT-S [64] 53.96 2.86M
CrossViT-S [80] 52.70 2.40M
LKCA-B 60.95 (Ours) 2.76M
T2T-B [77] 58.46 13.45M
CvT-B [69] 55.88 6.52M
MobileViTv2 [68] 58.28 8.17M
Twins_SVT-B [60] 49.24 9.04M
RegionViT-B [64] 57.83 12.39M
LKCA-L 63.43 (Ours) 12.65M
Table IV provides the comparison on Tiny-ImageNet. ViT and ViT-LKCA: The baseline ViT model achieves a Top-1 accuracy of 55.74% with 2.69M parameters and 174.25M Flops. In contrast, the ViT-LKCA model outperforms significantly with a Top-1 accuracy of 60.95% (an increase of 5.21 percentage points) and has the same 2.69M parameters and 170.81M Flops. Small-scale model (0.5M parameter magnitude): The ViT/0.5M model achieves a Top-1 accuracy of 46.53% with 0.48M parameters and 30.58M Flops. The ViT/0.5M-LKCA, incorporating LKCA enhancements, significantly improves accuracy to 50.59% (an increase of 4.06 percentage points) with 0.50M parameters and 31.34M Flops. Medium-scale models (1M, 2M, 4M parameter magnitudes): With ViT/4M-LKCA achieving a remarkable 6.14 percentage point improvement and reaching an accuracy of 62.30%. Large-scale model (8M parameter magnitude): For ViT and ViT-LKCA models with 8M parameters, ViT/8M-LKCA exhibits a significant improvement of 5.99 percentage points, achieving a Top-1 accuracy of 63.38%.
TABLE VI: Semantic Segmentation of Vanilla ViT and LKCA on ADE20K Dataset
Backbone Iteration Algorithm aAcc(%) mIoU(%) # Params
ViT-B 2000 UPerNet 40.90 22.28 0.144G
LKCA-B 2000 UPerNet 45.48 24.65 0.102G
ViT-B 4000 UPerNet 47.41 28.08 0.144G
LKCA-B 4000 UPerNet 52.67 30.00 0.102G
ViT-B 8000 UPerNet 53.43 31.73 0.144G
LKCA-B 8000 UPerNet 50.79 33.73 0.102G
ViT-B 16000 UPerNet 61.55 39.33 0.144G
LKCA-B 16000 UPerNet 65.49 40.60 0.102G
TABLE VII: Semantic Segmentation of Various ViT Variants and LKCA on ADE20K Dataset
Backbone Iteration Algorithm aAcc(%) mIoU(%) # Params Flops
Vit-S 10000 UPerNet 55.55 32.92 58M 67G
ResNet-50 10000 UPerNet 58.12 35.35 64M 238G
Twins_PCPVT-S 10000 UPerNet 52.56 29.64 53M 234G
Swin-S 10000 UPerNet 57.18 35.30 52M 231G
LKCA-S 10000 UPerNet 60.16 37.16 51M 66G
Vit-S 10000 FPN 56.56 35.84 37M 76G
Twins_PCPVT-S 10000 FPN 56.74 33.09 33M 50G
LKCA-S 10000 FPN 62.55 39.03 30M 78G
Vit-S 20000 UPerNet 58.51 36.61 58M 67G
ResNet-50 20000 UPerNet 62.76 37.84 64M 238G
Twins_PCPVT-S 20000 UPerNet 58.91 38.00 53M 234G
Swin-S 20000 UPerNet 58.00 37.57 52M 231G
LKCA-S 20000 UPerNet 64.37 40.09 51M 66G
Vit-S 20000 FPN 63.36 41.64 37M 76G
Twins_PCPVT-S 20000 FPN 58.51 36.61 33M 50G
LKCA-S 20000 FPN 67.21 41.81 30M 78G
@@
IV-C Comparison of ViT Variants and LKCA.
Next, we compared LKCA with other mainstream variants of Vision Transformer.
@@
From the information in the table I, in the context of small-scale models, Swin-T and CaiT-T perform similarly in terms of Top-1 accuracy, while MobileViTv2 lags slightly behind but possesses a significant advantage in computational efficiency. Compared to other small-scale models, LKCA-T achieves a higher Top-1 accuracy. Moving to large-scale models, it is a common trend that they generally outperform their small-scale counterparts. LKCA-B stands out by significantly surpassing ViT and MobileViTv2 in Top-1 accuracy, and it slightly outperforms CaiT-S and Swin-S. Overall, in CIFAR10 and CIFAR100, LKCA demonstrates higher Top-1 accuracy compared to other models. However, there is no clear advantage observed in the SVHN dataset.
Finally, we have validated the performance of various ViT variants and the LKCA model on the Tiny-ImageNet dataset, while also providing details on the models’ parameter counts and computational complexities in Table V.
@@
IV-D Semantic Segmentation Results